OpenStudy (anonymous):

Quadratic equations (3a-4)(2a-1)=0 ????

OpenStudy (anonymous):

hence \[a = \frac{ 4 }{ 3 }\] , \[a= \frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

could you explain how you got that

OpenStudy (anonymous):

when (3a-4)(2a-1) = 0 ... so, one of the equation in the bracket must be equal to zero. hence we got 3a-4 = 0 and 2a-1= 0 and then solve it...